Write a solutionset for y=3abs(7-2x)+5 when y<8.

Answer by stanbon(48558)

solutionset for y=3abs(7-2x)+5 when y<8
----------------
y=3|7-2x|+5 for y<8
8=3|7-2x|+5
|7-2x|=1
7-2x=1 or 7-2x=-1
2x=6 or 2x=8
x=3 orx=4
Solution:
3 Cheers,
Stan H.